OpenStudy (anonymous):

The sum of two numbers is 18. Three times the greater number exceeds four times the smaller number by 5. Find the numbers.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

i only know that there only 9 different ways to this

OpenStudy (texaschic101):

a + b = 18 --- a = 18 - b 3a = 4b + 5 now we sub 3(18 - b) = 4b + 5 54 - 3b = 4b + 5 54 - 5 = 4b + 3b 49 = 7b 49/7 = b 7 = b a + b = 18 a + 7 = 18 a = 18 - 7 a = 11 check.. 3a = 4b + 5 3(11) = 4(7) + 5 33 = 28 + 5 33 = 33 (correct) your 2 numbers are : 7 and 11